The resonant frequency of a closed air column is known to depend on the pressure and density of the air and length of the air column. How do these quantities relate?

1
Expert's answer
2022-02-14T14:28:48-0500

The resonant frequency F of a closed air column:

"f=\frac{V_s}{4l}"

We know that

"V_s=\sqrt\frac{E}{\rho}"

Now we can written as

"f=\frac{1}{4l}\times\sqrt{\frac{E}{\rho}}"

L = length of air colume